OpenStudy (anonymous):

Probability question: "Your neighbour has two children. You learn one of them is a boy, Sam. What is the probability that Sam's sibling is a brother?" (The answer is 1/3 ... please explain, thanks :)

OpenStudy (anonymous):

Could it be that possible cases: Brother / Sister / Undefined gender?

OpenStudy (anonymous):

No, it is only female or male.. and I think that conditional properties apply here; P ( Sam has a brother | Sam is male ) but I keep getting 1/2 regardless Please help is possible :)

OpenStudy (anonymous):

*if possible

OpenStudy (anonymous):

P ( Sam has a brother | Sam is male ) = P (Sam has brother n Sam is male) / P (Sam is male) P(Sam has brother n Sam is male) = (1/2)(1/2) = 1/4 P(Sam is male) = 1/2 P(condition) = 1/4 / 1/2 = 1/2 can't seem to get 1/3

OpenStudy (anonymous):

Same thing I did! But apparently the answer of 1/3 is correct :( according to my teacher as well

OpenStudy (raden):

the total number of outcomes = 3 {(M,M), (M,F),(F,M)} note : M=male, F=female the event that Sam's sibling is a brother = (M,M) so, the probility that Sam's sibling is a brother = 1/3

OpenStudy (anonymous):

Thank you so much! THat makes sense, because we have the premise that Sam is a male. :)

OpenStudy (raden):

yeah, this is an interesting problem :) You;re welcome